OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. @Mathematics

OpenStudy (anonymous):

y-5=2(x+1) y=2x+2+5 y=2x+7